An element with mass 730 grams decays by 28.8% per minute. How much of the element is remaining after 10 minutes, to the nearest 10th of a gram?



1
Expert's answer
2022-03-07T14:51:03-0500

Decay


"m(t)=m(0)e^{kt}"

Given "m(0)=730\ g"


"m(1)=m(0)e^{k(1)}=(1-0.288)m(0)"

"k=\ln0.712"


"m(t)=730e^{(\ln0.712)t}=730(0.712)^t"

"m(10)=730(0.712)^{10}"

"m(10)=24.4\ g"